Prove that if 10 points are placed in a 3cm by 3cm squre board than two points must be at least √2

cm apart.


1
Expert's answer
2021-11-23T18:04:24-0500

A 3cm by 3cm squre board are divided by nine 1cm by 1cm squares. Since there are 10 points, by the Pigeonhole Principle there is a 1cm by 1cm square that contains at least two points. Since the maximal distance between any two points in the square is equal to the length of a diagonal which is equal to "\sqrt{1^2+1^2}=\sqrt{2}" for the 1cm by 1cm square, we conclude that two points must be at least "\sqrt2" cm apart.